J is between H and K. If HJ = 6x - 5, JK = 4x -6, and KH = 129, find the value of x.



1
Expert's answer
2020-10-28T16:54:47-0400

HK = KH = 129

HJ = 6x - 5

JK = 4x - 6


HK = HJ + JK

HK = 6x - 5 + 4x - 6

HK = 6x + 4x - 5 - 6

HK = 10x - 11


Since HK = 129


Therefore,

10x - 11 = 129

10x = 129 + 11

10x = 140

Divide both sides by 10

"\frac{10x}{10} = \frac{140}{10}"


x = 14